Startup Night 5 - 20170906 UT

Observers: JHill, BRothberg, A.Cardwelll (Tucson)
Telescope Operator: G.Bechetti (LBTO)
AO Support: D.Miller
Telescope Support: P.Hartley (LBTO)
Software Support: CBiddick, KSummers (Tucson)
Instrument Support: ESolheid (LBTO)
Engineering Support: McMahon

Plan

  • LBC Twilights at Start
  • LBC Photometric and Spectro-Photometric Standard Fields
  • Field Aberrations SX and DX
  • AGW 1 Transform
  • AGW 2 Confirmation of new Transform
  • LUCI Science OB tests
  • - Longslit Acquisition
  • - Non-Sidereal SL test
  • - Non-Sidereal AO (functional check only)
  • AGW4 (Pepsi)
  • LBTI M3 Alignment Check

Log

UT 01:40 - Started LBCs OK, LUCI Software (Observer GUI) in bad state - called T. Sargent

UT 01:51 - Blue Trackers issue on LBC, LBCKILL/RESTART ISSUED - attempting Uspec and N972N20 first

UT 01:52 - LUCI Observer GUI/java software fixed - issue was software binaries were not updated on the RM525 and OBS machines - all working now

UT 01:53 - Sky Brightness 10:49 - Uspec and N972 Started PA=0
  • 015210 015210 started exposures
  • Uspec saturated, F972 finished
  • Running F972 PA=180, blue channel turned off (saturated)
  • Sky Brightness 13.78
  • Scaling Factor 1:10 for Uspec and N972 - PA 180 counts now about equal between the two
UT 02:11 - Sky Brightness 15.18 - Uspec and i-Sloan PA=180, then PA=0

UT 02:34 - odd occurrence in the LBC GUI - showing completed more exposures than part of script (i.e. 7/5 completed)

UT 02:41 - moved to standard SA113 - dohybrid cannot detect

UT 02:48 - moving to GD248 - too low

UT 02:50 - moving to G24-9 (specstd) - will collimate then do co-pointing then run script - Keck notes this is an eclipsing binary - so best not to use - still worth for collimating

UT 03:00 - moving to LDS749B (specstd - this one looks good) - running DOFPIA as a check, then will co-point

UT 03:12 - starting exposures for LDS749B Airmass 1.530, DIMM 0.83"

UT 03:51 - completed SpecStd - reconfiguring for LUCI 1 & 2

UT 03:58 - will run field aberration test using Doug M. scripts - field is BS 9177 (Ra 18hr 39m +49deg 05min) - list is "BS9177_new.list"

UT 04:19 - SX and DX AdSecs need a restart

UT 04:24 - J. Hill updating with new pointing model from D. Summer, restart PCS

UT 04:49 - Sent BS9177_new.list - note: the default search is for the word "field" to be in the IDL popup

UT 05:02 - failedon GS13, restarted with start=3 (SX side)
  • IRCL: preset, BS9177, MODE=ACTIVE, GS=0, PA=89.0897 (iif_preset)

    => left preset complete
    261 nm rms error in image left_wfscimage000013.fits
    => Collimated in 1 iteration

    IRCL: preset, BS9177, MODE=ACTIVE, GS=13, PA=169.167 (iif_preset)

    ===> Error in irc_presetTelescope
    err = Resmsg[0] = PresetTelescope result status: Error
    err = Resmsg[1] = Stop{FLAO}[214]
    err = Resmsg[2] = guiding was not stable within 6 cycles or lost GS

  • Repeat of this worked OK
  • GS14 on SX: might have picked the wrong star
  • GS33 DX - wrong star picked, its a binary
  • GS30 SX - vanished - didn't find star resending UT 05:44
  • restarted SX side with start=11 UT 05:45, GS30 on SX failed to find star
  • restarted SX side with start=12 UT 05:48
  • SX didn't find GS0 - after restart
  • Homing AGW1 probe and restarting GCS left
  • restarting SX at start=11 (GS30) to see if restart of GCS fixed issue
  • SX still failed to find even GS0 - which is odd as it started OK with GS0
  • DX is finished at UT 05:52
  • Moving to pointing star to check on SX - only completed 7 or 8 so worth checking pointing
  • restarting SX at start=11 ==> these scripts are not binocular encoded - OSA has to manually take us to star
  • going back to field and finishing last 3 on SX, then will take 3 more on DX
  • SX seems happy with GS30
  • SX GS33 is a double - ignore it
  • DX at GS0 after GS13 had a jumpy guidestar acquisition
  • SX finished
  • GS13 last one on DX

UT 05:15 - Moon seems to illuminate some thin wispy clouds in the east

UT 05:19 - DIMM reports 0.75" about the same on GCS

UT 06:17 - moving to Vy2-2 for test acquisition and LS data Binocular mode

UT 06:22 - guidestar acquired on both sides OK (DX was close to edge)

UT 06:30 - DX active optics wrinkly - psf blows up and then settles down

  • 1pixel -0.12" offset for DX - SX looks good
  • DX looks like it overshot,but its bounced in multiple reads in the GEIRS display - settled down to center of slit
  • Starting Science exposure - 60sec might be too bright for Pa Beta line (might be saturated)
  • Sync error GEIRS on LUCI-2 Local 23:42 - could not add FITS properties, error commiting setup to readout manager
  • manually took image again and it saved OK
  • Loaded HKspec script, skipped to spectroscopic config line to take data (Airmass 1.49)
  • Br Gamma might be saturated or close to it in 30sec
  • Negative countdown on GEIRS2 - error again, pressed reset on LUCI2 Readout manager - looks GEIRS2 froze, manually resubmitted (Local 23:50 - same error could not add FITS properties)
ID: 13930515
Software Timestamp: 2017-09-06 06:50:20.4630
Time Service Client: sync
Type: ERROR (12)
Level: HIGH (2)
Usage: Luci TWO (2)
Host: luci.luci.lbto.org (137)
Program: TWO_GEIRSServer (11)
Address: RMIGEIRSServiceImpl.java:805
long Address: de.rub.astro.lucifer.control.readout .RMIGEIRSServiceImpl .checkOkAnswer :805
Content: GEIRS received response with error: ERROR 1 Command 'sync read' returned errorcode = 1: (E_error=1) general error: 'error: ERROR read (E_ptimeout=21) timeout on OPTPCI interface' sh: geirs_dropcaches: command not found .
Attachment:  
Details: Transmission

ID: 13930519
Software Timestamp: 2017-09-06 06:50:20.7890
Time Service Client: sync
Type: ERROR (12)
Level: HIGH (2)
Usage: Luci TWO (2)
Host: luci.luci.lbto.org (137)
Program: TWO_GEIRSServer (11)
Address: RMIGEIRSServiceImpl.java:1145
long Address: de.rub.astro.lucifer.control.readout .RMIGEIRSServiceImpl .updateFITSHeader :1145
Content: Could not add FITS properties ("activatable server "TWO_GEIRSServer" at registry "rmi://localhost:60002/(127.0.0.1):60002"").
Attachment: java.lang.NullPointerException
Details: Transmission

ID: 13930521
Software Timestamp: 2017-09-06 06:50:21.9970
Time Service Client: sync
Type: ERROR (12)
Level: HIGH (2)
Usage: Luci TWO (2)
Host: luci.luci.lbto.org (137)
Program: TWO_GEIRSServer (11)
Address: RMIGEIRSServiceImpl.java:805
long Address: de.rub.astro.lucifer.control.readout .RMIGEIRSServiceImpl .checkOkAnswer :805
Content: GEIRS received response with error: ERROR 1 Command 'sync save' returned errorcode = 1: (E_error=1) general error: 'error: ERROR save error: framebuffer is empty (not yet read?)' .
Attachment:  
Details: Transmission

  • resubmitted worked OK
  • Telluric for Vy2-2 - HD188920
  • Started Telluric spectroscopy UT 07:02 (zJspec) Airmass 1.36
  • Started Telluric spectroscopy UT 07:09 (HKspec) Airmass 1.38
  • Tilt Error LUCI-1 unable to stablize at 1.93 micron UT 07:10
  • 13933125 2017-09-06 07:10:07.5930 ERROR *** System Scheduler Error committing setup to instrument manager at "rmi://luci.luci.lbto.org:60001/ONE_InstrumentManager"!
    13933124 2017-09-06 07:10:07.5590 SYSTEM *** Luci ONE ONE_GratingUnit grating unit at position 2, order 1, wavelength 1.9299999475479126, tilt voltage 0.28474700450897217, loop closed, tilt ...
    13933123 2017-09-06 07:10:07.5330 ERROR *** Luci ONE ONE_InstrumentManager error while committing setup GRATING
    13933122 2017-09-06 07:10:07.5320 ERROR *** Luci ONE ONE_InstrumentManager Error tilting the grating unit to 1.9299999475479126 ?m at "rmi://localhost:60001/ONE_GratingUnit"!
  • Manually changed LUCI-1 grating to 1.93 and manually took exposure
  • Vy2-2 and Telluric are #0002-0024 for LUCI1 & LUCI2

UT 07:23 - Attempting non-sidereal SL LUCI imaging (binocular) of target "7 IRIS" from A. Conrad. First will do a pointing check then test NSIGUI. Main belt asteroid, 4th brightest in belt.

UT 07:26 - error on preset for pointing check, had accidentally left in non-sidereal hijack when Geno B. did a pointing check.

UT 07:29 - pointing good, hijack set, preset sent - failed

UT 07:34 - EXP #0025 on LUCI 1 & 2 is 2.8sec X 22 DIT H-band image - useless -guiders are on-axis

UT 07:40 - remade script with Track mode but forgot to hijack preset

UT 07:41 - trying again - preset failed on one side, didn't park probe on SX

UT 07:42 - trying again - failed

UT 07:48 - trying again in track mode with hijack - however, this ephemeris is Osiris-Rex not 7 IRIS (NSIGUI problem won't load new stuff)
  • #0026 on both sides
  • Stopped on "RUNNING" LUCI-1 side, but exposure is saved, skipped to next line and clicked GO
  • Finished H-band and 1 J-band don't see anything
  • John H. has found the issue and correcting the ephemeris
  • Found field coords close to 08:15 UT picked guidestar sending guiding with NS @UT08:11
  • preset failed - reloaded from NSIGUI
  • still failed - could be a punctuation issue (7IRIS has a ";" in it - causing us to move far off target)
  • DIMM says seeing is 1.45"
  • Still failed - unique identifier for Horizons has ";" even though we renamed ephemeris file this still fails
  • changed to "=" still fails, we are restarting IIF
  • restart DD and restart telescope service on LUCI because IIF restarted
  • PCS failed
  • Restarting PCS
  • Right BFG rotator also turned off (how is this related?)
  • Captain Needa's lucky day?
  • Barry reloaded the ephemeris, said "Watch this!" and sent the hijacked preset
  • Moving somewhere different (facsum flashed yellow - wasn't supposed to do that)!
  • Sorry Captain Needa....we hit our target!
  • nevermind we hit the AdSec limits
  • resending Track Mode version of script (DIMM shows 1.75")
  • not sure we see anything
  • taking 9 H-band exposures, 6 J-band exposures, 3 z-band NS track mode only - #0033-0049

UT 09:02 - Going to try making an AO-script for NS. Andrew C. will make script with bright star - will need to use the ARGOS version of the OT

UT 09:20 - Script made, waiting on AO startup

UT 09:29 - John H. running tests on hijacking and pointing

UT 09:36 - sending script, LUCIs configured, waiting for AO loops to close, DIMM = 1". The hijacked ACE_TRACK preset succeeded. The AORef magnitude is the one from the script and not the ephemeris file.

UT 09:41 - RunAO failed - taking test image on LUCI-2 2.8sec

UT 09:47 - taking another 2.8sec x 22DIT test image on LUCI-2

UT 09:56 - resending preset from script. BUT WE FORGOT TO HIJACK. This was successful although Doug had to use intervention mode on the SX camera lens centering.

UT 10:16 - resending preset from script - really NS this time. This was successful although Doug had to use intervention mode on the SX camera lens centering. Dither pattern of 10 arcsec seems to big for N30 field.

UT 10:30 - some delay at a Pause in the script...... Pause was left in from previous attempt, but another pause occurred after this unrelated to an inserted pause. Last exposures: L1 #0057; L2 #0059

UT 10:51 - Pointing and collimation model tested again - seeing highly variable from 1.8" to 1"

UT 11:14 - Reconfigured for LBCs

UT 11:31 - moving to co-pointing for G191B2b - will collimate and co-point

UT 11:42 - sending G191B2b OB

UT 11:58 - completed OB

UT 11:59 - moving to start skyflats

UT 12:00 - running r-sloan r-sloan test PA=0

UT 12:12 - starting r-sloan, r-sloan, PA=0 scaletime = 4sec for both
  • 4k counts first blue, 4.7k counts red
  • 2nd exposures ~5k counts, ended at 6k counts
  • running again, 4sec both PA=0, 2nd exposure 7k counts, ends at 12k counts

UT 12:22 0 r-sloan, r-sloan, PA=180 - scaletime 1.5sec
  • starting at 8k,finished 13k-15k

UT 12:26 - g-sloan, I-Bessel PA=180 scaletime=1sec -already hitting 35k -too high -stopping OB finished at 4/10

UT 12:28 - B-Bessel, V-Bessel PA=180 exptime 1sec - 18k already at start - will let this run

UT 12:33 - done with skyflats turning off LBCs and closing enclosure -will set up for LUCI 1&2 internal calibrations

UT 12:43 - will test Jenny P. fraternal G200 HK,zJ longslit scripts
  • got an error - empty wavelength, error committing, grating central wavelength not specificed - they are specified in the OT Beta
  • Going to mono - LUCI-1 zJ flats scripts LUCI1 #0058-0067
  • Will do LUCI-2 in parallel, but by hand - starting with zJ once mask is in place LUCI2 #0058-0067 (Halo1)
  • LUCI-1 HKflats #0068-0077 -script
  • LUCI-2 HKflats #0068-0079 -by hand (Halo2)
  • LUCI-1 HK Arcs #078-80 -script - ERROR - script doesn't appear to take darks - will take 3 darks of 20sec manually #081-083
  • LUCI-2 HK Arcs #080-81 (Ne), #082-083 (Ar), #084-085 (Xe) -20sec each
  • LUCI-1 zJ Arcs #084-085 - take darks manually too they are missing (60 sec x2) - IGNORE 084-085 forgot to set to zJ - HKspec shows too much thermal noise
  • #086-087 Arcs, 088-089 60sec Darks
  • LUCI-2 zJ Arcs #086-087 (Ne), #088-089 (Ar) -60sec each
  • Finished Calibrations for on-sky Vy2-2 data UT 13:26

Summary

  1. Field Aberration test run (SX and DX) - used updated pointing and collimation model from Doug S.
  2. LUCI G200 zJ and HKspec longslit acquisition and collection of data and telluric run. IT 6050 (L1 grating stabilization) and IT 6555 (negative GEIRS2 countdown with error and restart of GEIRS2) noted.
  3. Non-sidereal LUCI-Binocular AO tested and apparently works.
  4. Non-sidereal LUCI SL (track only) tested - some issues regarding Horizons and nomenclature. Some difficulty with LUCI SL non-sidereal guiding - could be related to Horizons issue, was difficult to get guidestar.
  5. LBC spectrophotometric standard observed at start of night (LDS749B, G191B2b)
  6. Pointing and collimation checks for new model - earlier in night large travel from +61 dec to +09 dec was successful, guidestars on both sides acquired
  7. LBC Twilight flats at PA=0 and 180: LBC-B: Uspec, r-sloan ; LBC-R: r-sloan, sloan i, N972. PA=180 only for B-Bessel and V-Bessel but counts might be too high (counts 18k-37k)
  8. G200 zJ and HK calibrations (0.5" LS) - tested fraternal OT Beta scripts (did not work), verification that previous exposure times are still good for both LUCIs (NOTE: LUCI-1 script doesn't take darks)
 
Topic revision: r3 - 09 Sep 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback